Example Presentation Created with the Beamer Package

Till Tantau

June 10, 2009

Till Tantau  Example Presentation Created with the Beamer Package

Features of the Beamer Class

▶ Normal LaTeX class. ▶ Easy overlays. ▶ No external programs needed.

▶ Normal LaTeX class. ▶ Easy overlays. ▶ No external programs needed.

▶ Normal LaTeX class. ▶ Easy overlays. ▶ No external programs needed.